### v2

# RAGONITE: Iterative Retrieval on Induced Databases and Verbalized RDF for Conversational QA over KGs with RAG

Abstract questions like ‘Innovative highlights in x7?’ require NL understanding and common sense reasoning to map the intent to equipments and accessories of the BMW X7. We address such questions by verbalizing KGs [Og22], so that LLMs can reason over them like NL text. We again use the KG in Turtle format, where all facts with a specific entity as subject are grouped together. The facts in each Turtle “capsule” are then verbalized into NL passages using simple rules: (i) all schema-related prefixes are stripped to extract the raw content; (ii) the subject string, predicate string and object string are concatenated to form a sentence; (iii) “is” is added before type predicates, and “has” before other predicates; and (iv) The reverse formulation of the fact is also added to the passage to later help the generator LLM answer questions when the contained information is requested in complementary ways (What is BMW 120 Sport’s engine performance? and Which BMW engines have a performance of 125 kW?). Explicitly adding reversed facts is helpful for smaller LLMs.
A toy RDF Turtle capsule looks like: <engine/bmw-120-sport> a ns1:EngineSpecification ; ns1:enginePerformance "125 kW"; ns1:fuelType <fuel-type/gasoline> . This is verbalized into an NL passage as follows: “BMW 120 Sport is Engine Specification. BMW 120 Sport has engine performance 125 kW. 125 kW is engine performance of BMW 120 Sport. BMW 120 Sport has fuel type gasoline. Gasoline is fuel type of BMW 120 Sport.”
2.3 Iterative retrieval
Abbildung 3: Illustrating iterative retrieval with SQL and text search tools via LLM calls in RAGONITE.
It is possible that results from either branch of RAGONITE are not of sufficient quality. For example, the SQL tool might generate a partial query or the text results may be incomplete or contain only irrelevant information. For such cases, the LLM can request additional rounds (set to three, but configurable) of retrieval from either branch. Retrieval results, and possibly error messages from one round are included in the prompt in the next round. Sending such errors back to the retrievers gives them the opportunity to correct small mistakes, like resolving an ambiguous column name (for SQL), and guiding subsequent searches towards more relevant evidence (for text search). We further force the LLM to use both branches at least once to ensure as comprehensive a response as possible. This iterative workflow with tools is shown in Fig. 3.
2.4 Branch integration
At the end of iterative retrieval, a second LLM agent generates an answer using both the SQL outputs and the top passages being inserted into its prompt. This way we do not have to decide which branch to prefer, neither at question-time nor after obtaining an answer from each mode: with both results at its disposal, the LLM can decide how to integrate them. The integration of branches and judgment over individual results is thus automated via the LLM and are not hard-coded into the system.
2.5 Open LLM support
The default LLM in RAGONITE is GPT-4o with API access. A desideratum of agentic LLM systems is to ensure data security via local, on-premise LLM hosting. Along these lines, RAGONITE supports local deployment of open LLMs such as Llama-3.3. For our experiments, we hosted the 70B version with 4 bit-quantization using ollama [^3] https://ollama.com/library/llama3.3.
2.6 Heterogeneous QA
Since one of our retrievers runs on NL verbalizations of the KG, it is straighforward to insert additional text contents (for example, Web documents with supplementary information to what is contained in our KG) to our backend knowledge repository. In one version of the demo, we incorporated about 400 English passages from the BMW website [^4] https://www.bmw.co.uk/en/index.html to increase the scope of answerable questions: RAGONITE then becomes capable of heterogeneous question answering with a mix of a KG and a text collection as its knowledge sources.
3 Demonstration walkthrough
Abbildung 4: Screenshot of RAGONITE. Colored boxes are not part of the UI. Text readable on zoom-in.
Overview. The UI in RAGONITE (Fig. 4) is divided vertically into four panels. Left to right, (i) the first panel stores previous chats; (ii) the main chat panel takes questions as input and displayes system answers; (iii) the third panel shows answer derivation steps; and (iv) the last strip helps switch between different RAGONITE configurations.
Walkthrough. We go through only the main workflow here: everything else can be explored by a user during an interactive demo session. Offline, the backend KG is preprocessed to derive the equivalent DB and the text corpus. At runtime, in Step 1, users enter their conversational question (potentially intent-implicit) in the input box provided. Suppose we are at the second conversation turn now (Sounds like these are sleek sedans. Is an average X1 taller than a Coupe?). In Step 2a, this question is passed on to the intent-explicit SQL-formulating LLM call (Sec. 2.1). In this case, the LM decides that results of one iteration (that only retrieved heights of BMW X1 variants) was not enough to satisfy the intent in this complex question (Sec. 2.3), and so it formulates a second query (this time fetching the heights of Coupe models). At this moment, the iterator stops SQL querying and invokes the text retriever (Step 2b, Sec. 2.2) via an intent-explicit NL question (BMW X1 average height compared to BMW Coupe Models). Note that Steps 2a and 2b can happen in an arbitrary order. A sample verbalization of a Turtle capsule can be seen in the screenshot below the corresponding blue box (for the entity “bmw x1 xdrive 25e sport”, we lowercase all text for simplicity), along with the retrieval score assigned by the vector search model (0.01700170.0170.017). The iterator decides that contents of one round of text retrieval has enough information pertinent to the question, and by now both tools have been explored, so it hands over these results to the answer generator (Sec. 2.4) This LLM agent crafts the final response shown to the user (Step 3). Through explicit prompting, we make the LLM inject source citations in its answer for scrutability. In this case, sources [1-2] refer to SQL results and [3-7] to verbalizations (totaling $10$ sources, where $10$ is configurable). Prompts for all LLM calls can be examined by an user via the answer derivation panel.
Backend. The core of RAGonite’s backend handles query formulation, SQL execution, text retrieval, iteration/tool selection, and answer generation. The backend also has a stateful layer that stores conversations in a SQLite DB and provides a REST API to the frontend with FastAPI. Core dependencies include the vector database (ChromaDB), a template engine used for prompts (Jinja), and LLM libraries (OpenAI and Ollama). ChromaDB is used as our vector DB for storing RDF verbalizations. We also use ChromaDB’s in-built vector search. For all of our experiments, we ran RAGONITE on a GPU server (4x48GB NVIDIA Ada 6000 RTX, 512 GB RAM, 64 virtual cores) which is shared with other researchers. A single GPU was dedicated to running Llama-3.3-70B. If the LLM used for text inference is hosted in the cloud, for example, by using an OpenAI model, the remaining resource requirements are minimal and the RAGonite instance can also run on a standard laptop.
Frontend. We created a single-page React app, intentionally avoiding further dependencies to preclude the need for a build process. All API calls are handled by the frontend.
4 Evaluation
Setup. Our BMW KG (scraped from bmw.co.uk) has $3442$ facts, $466$ unique entities, $27$ predicates, $7$ types, and $1295$ literals. So we get $7$ tables and $466$ passages via DB induction and RDF verbalization. The authors generated $6$ conversations with $5$ turns each, and test $4$ GPT-4o-based configurations with them: (i) SPARQL-only, (ii) SQL-only, (iii) verbalization-only, and (iv) SQL+verbalizations. Questions were in $3$ categories ($10$ in each): (i) lookup intents (cost of 530em sport saloon?), (ii) complex intents (charging time of 225e Active Tourer less than avg over all bmws?), and (iii) abstract intents (luxury features in x5?). Answers were marked correct/incorrect (SQL/SPARQL query must also be correct).
Results. The two-pronged iterative approach did the best ($28/30$ questions correct). SQL-only got $18$ (failure to handle abstract intents) and verbalization-only got $24$ cases right (failure to handle complex math). A baseline where GPT-4o generated SPARQL queries, only got $4$ answers correct. A particular vulnerability of SQL and SPARQL was incorrect entity linking for ad hoc mentions (like “gran coupe sport 220 i m” in questions, while the exact entity label is bmw-220i-m-sport-gran-coupé), and this is where verbalizations are particularly helpful. This also proved that LLMs are still much weaker at SPARQL generation than SQL (presumably due to lower training data), especially when it comes to complex or abstract intents in ConvQA.
Runtimes. We recorded stepwise runtimes for each of our $30$ benchmark questions using our default pipeline with GPT-4o. Including question completion, the entire retrieval took 3.93639363.9363.936 seconds, and answering took 2.40924092.4092.409 seconds. Zooming in on the retrieval step, tool selection took 1.04710471.0471.047 seconds, SQL execution took 0.09500950.0950.095 seconds, and vector search took 0.53905390.5390.539 seconds. The entire RAG pipeline thus took 6.34863486.3486.348 seconds per question, on average. RAGONITE thus operates with interactive response times.
All public artifacts (like code, data, and demo screenshots) related to this work can be found at https://github.com/Fraunhofer-IIS/RAGonite.
5 Conclusions and future work
We showcase RAGONITE, a transparent RAG pipeline that adopts a novel alternative approach to ConvQA over KGs by iterative retrieval and merging results of structured and unstructured querying. While Text2SPARQL has been investigated in the past, we adopt a novel alternative of exploiting Text2SQL, where LLMs do much better, over a DB that is programmatically induced from the KG. By converting a KG to its verbalized form that is indexed as a vector DB, we show that the information in knowledge graphs can be harnessed for more abstract intents as well. RAGONITE contains an agentic workflow: key future work would enhance this further by incorporating reflection mechanisms that enable each module to critique and improve its output, without compromising efficiency. Another direction would be to fine-tune specific pipeline components for a RAG setting with synthetic data.
Acknowledgements
This research was funded by the German Federal Ministry for Economic Affairs and Climate Action (BMWK) through the project OpenGPT-X (project no. 68GX21007D). We thank Lucas Weber from the NLP team at Fraunhofer IIS for useful comments on an initial version of this manuscript. We also thank the rest of the team members for useful inputs at various stages of this work, intensively testing our demo, and suggesting benchmark questions.
Literatur
  • [Ac23] Achiam, Josh; Adler, Steven; Agarwal, Sandhini; Ahmad, Lama; Akkaya, Ilge; Aleman, Florencia Leoni; Almeida, Diogo; Altenschmidt, Janko; Altman, Sam; Anadkat, Shyamal; Avila, Red; Babuschkin, Igor; Balaji, Suchir; Balcom, Valerie; Baltescu, Paul; Bao, Haiming; Bavarian, Mohammad; Belgum, Jeff; Bello, Irwan; Berdine, Jake; Bernadett-Shapiro, Gabriel; Berner, Christopher; Bogdonoff, Lenny; Boiko, Oleg; Boyd, Madelaine; Brakman, Anna-Luisa; Brockman, Greg; Brooks, Tim; Brundage, Miles; Button, Kevin; Cai, Trevor et al.: GPT-4 technical report. arXiv, 2023.
  • [Al24] Ali, Mehdi; Fromm, Michael; Thellmann, Klaudia; Ebert, Jan; Weber, Alexander Arno; Rutmann, Richard; Jain, Charvi; Lübbering, Max; Steinigen, Daniel; Leveling, Johannes; Klug, Katrin; Buschhoff, Jasper Schulze; Jurkschat, Lena; Abdelwahab, Hammam; Stein, Benny Jörg; Sylla, Karl-Heinz; Denisov, Pavel; Brandizi, Nicolo’; Saleem, Qasid; Bhowmick, Anirban; Helmer, Lennard; John, Chelsea; Suarez, Pedro Ortiz; Ostendorff, Malte; Jude, Alex; Manjunath, Lalith; Weinbach, Samuel; Penke, Carolin; Filatov, Oleg; Asaadi, Shima; Barth, Fabio; Sifa, Rafet; Küch, Fabian; Herten, Andreas; Jäkel, René; Rehm, Georg; Kesselheim, Stefan; Köhler, Joachim; Flores-Herr, Nicolas: Teuken-7B-Base and Teuken-7B-Instruct: Towards European LLMs. arXiv, 2024.
  • [Ba20] Bakhshi, Mahdi; Nematbakhsh, Mohammadali; Mohsenzadeh, Mehran; Rahmani, Amir: Data-driven construction of SPARQL queries by approximate question graph alignment in question answering over knowledge graphs. Expert Systems with Applications, 146, 2020.
  • [Ch19] Christmann, Philipp; Saha Roy, Rishiraj; Abujabal, Abdalghani; Singh, Jyotsna; Weikum, Gerhard: Look before you hop: Conversational question answering over knowledge graphs using judicious context expansion. In: CIKM. 2019.
  • [Di20] Diefenbach, Dennis; Giménez-García, José; Both, Andreas; Singh, Kamal; Maret, Pierre: QAnswer KG: designing a portable question answering system over RDF data. In: ESWC. 2020.
  • [Ga17] Galkin, Mikhail; Auer, Sören; Vidal, Maria-Esther; Scerri, Simon: Enterprise knowledge graphs: A semantic approach for knowledge management in the next generation of enterprise information systems. In: ICEIS. 2017.
  • [Go17] Gomez-Perez, Jose Manuel; Pan, Jeff Z; Vetere, Guido; Wu, Honghan: Enterprise knowledge graph: An introduction. In: Exploiting linked data and knowledge graphs in large organisations. Springer, 2017.
  • [Ji24] Jiang, Albert Q; Sablayrolles, Alexandre; Roux, Antoine; Mensch, Arthur; Savary, Blanche; Bamford, Chris; Chaplot, Devendra Singh; Casas, Diego de las; Hanna, Emma Bou; Bressand, Florian; Lengyel, Gianna; Bour, Guillaume Bour; Lample, Guillaume; Lavaud, Lélio Renard; Saulnier, Lucile; Lachaux, Marie-Anne; Stock, Pierre; Subramanian, Sandeep; Yang, Sophia; Antoniak, Szymon; Le Scao, Teven; Gervet, Théophile; Lavril, Thibaut; Wang, Thomas; Lacroix, Timothée; El Sayed, William: Mixtral of experts. arXiv, 2024.
  • [Ka20] Karpukhin, Vladimir; Oguz, Barlas; Min, Sewon; Lewis, Patrick; Wu, Ledell; Edunov, Sergey; Chen, Danqi; Yih, Wen-tau: Dense Passage Retrieval for Open-Domain Question Answering. In: EMNLP. 2020.
  • [Ka21] Kacupaj, Endri; Plepi, Joan; Singh, Kuldeep; Thakkar, Harsh; Lehmann, Jens; Maleshkova, Maria: Conversational Question Answering over Knowledge Graphs with Transformer and Graph Attention Networks. In: EACL. 2021.
  • [KSRW21] Kaiser, Magdalena; Saha Roy, Rishiraj; Weikum, Gerhard: Reinforcement learning from reformulations in conversational question answering over knowledge graphs. In: SIGIR. 2021.
  • [Li24] Li, Jinyang; Hui, Binyuan; Qu, Ge; Yang, Jiaxi; Li, Binhua; Li, Bowen; Wang, Bailin; Qin, Bowen; Geng, Ruiying; Huo, Nan; Zhou, Xuanhe; Ma, Chenhao; Li, Guoliang; Chang, Kevin C.C.; Huang, Fei; Cheng, Reynold; Li, Yongbin: Can LLM already serve as a database interface? A Big bench for large-scale database grounded Text-to-SQLs. In: NeurIPS. 2024.
  • [Me24] Meyer, Lars-Peter; Frey, Johannes; Brei, Felix; Arndt, Natanael: Assessing SPARQL capabilities of Large Language Models. In: NLP4KGC. 2024.
  • [Og22] Oguz, Barlas; Chen, Xilun; Karpukhin, Vladimir; Peshterliev, Stan; Okhonko, Dmytro; Schlichtkrull, Michael; Gupta, Sonal; Mehdad, Yashar; Yih, Scott: UniK-QA: Unified Representations of Structured and Unstructured Knowledge for Open-Domain Question Answering. In: NAACL. 2022.
  • [PB24] Perevalov, Aleksandr; Both, Andreas: Towards LLM-driven Natural Language Generation based on SPARQL Queries and RDF Knowledge Graphs. In: Text2KG. 2024.
  • [RA22] Roy, Rishiraj Saha; Anand, Avishek: Question Answering for the Curated Web Tasks and Methods in QA over Knowledge Bases and Text Collections. Springer, 2022.
  • [Ra24] Rangel, Julio C; de Farias, Tarcisio Mendes; Sima, Ana Claudia; Kobayashi, Norio: SPARQL Generation: An analysis on fine-tuning OpenLLaMA for Question Answering over a Life Science Knowledge Graph. arXiv, 2024.
  • [Sa18] Saha, Amrita; Pahuja, Vardaan; Khapra, Mitesh; Sankaranarayanan, Karthik; Chandar, Sarath: Complex sequential question answering: Towards learning to converse over linked question answer pairs with a knowledge graph. In: AAAI. 2018.
  • [So17] Song, Dezhao; Schilder, Frank; Hertz, Shai; Saltini, Giuseppe; Smiley, Charese; Nivarthi, Phani; Hazai, Oren; Landau, Dudi; Zaharkin, Mike; Zielund, Tom; Molina-Salgado, Hugo; Brew, Chris; Bennett, Dan: Building and querying an enterprise knowledge graph. IEEE Transactions on Services Computing, 12(3), 2017.
  • [Su20] Sun, Yawei; Zhang, Lingling; Cheng, Gong; Qu, Yuzhong: SPARQA: skeleton-based semantic parsing for complex questions over knowledge bases. In: AAAI. 2020.
  • [To23] Touvron, Hugo; Lavril, Thibaut; Izacard, Gautier; Martinet, Xavier; Lachaux, Marie-Anne; Lacroix, Timothée; Rozière, Baptiste; Goyal, Naman; Hambro, Eric; Azhar, Faisal; Rodriguez, Aurelien; Joulin, Armand; Grave, Edouard; Lample, Guillaume: LLaMA: Open and efficient foundation language models. arXiv, 2023.
  • [Va21] Vakulenko, Svitlana; Longpre, Shayne; Tu, Zhucheng; Anantha, Raviteja: Question rewriting for conversational question answering. In: WSDM. 2021.
  • [Ya23] Yang, Shuangtao; Teng, Mao; Dong, Xiaozheng; Bo, Fu: LLM-based SPARQL generation with selected schema from large scale knowledge base. In: CCKS. 2023.